OpenStudy (purple_pink):

find 2 ratios that are equivalent to the given ratio 72 : 36

OpenStudy (purple_pink):

@TheProfessor35

OpenStudy (purple_pink):

brb in 1 min

OpenStudy (purple_pink):

i gtg do something

OpenStudy (theprofessor35):

72/2 = 36 36/2 = 18 36:18

OpenStudy (theprofessor35):

One more? 72/36 = 2 36/36 = 1 2:1

OpenStudy (purple_pink):

m back

OpenStudy (theprofessor35):

I explained above ^

OpenStudy (purple_pink):

so you can divide?

OpenStudy (theprofessor35):

Yes, as long as you apply the SAME number to BOTH numbers. If you choose divide by 2, do it to both sides.

OpenStudy (purple_pink):

thanks 2 more